SW&IT Meeting Minutes 2019-04-22
Attendees: Andrew, Matthieu, Petr, Stephen, Yang
Report auto-generated Apr 22 6.45am, Do not
edit this table, use entries in smartsheets.
| LR20190408 - Bec, Matthieu
| LR20190408 - Colson, Andrew
| LR20190408 - Edmeades, David
| LR20190408 - Sola, Igor
| LR20190408 - Kubanek, Petr
| LR20190408 - Hooper, Stephen
| LR20190408 - Zhang, Yang
| LR20190408 - Zhang, Xianyu
News from management
- CH panel interview Tue Apr 23
Previous Action Items
- (last week) Qt webinar, Qt will support https://webassembly.org sometime soon (mid-2019?)
- OSS telescope tests
- TCSConfigFile fix/tests
- test target in Makefiles
- MCSPU new binaries, shall fix core dumps
- contacted M.Lesser about AzCam simulator; as none is available, I will code one based on oac/hkserver (one is really needed for GCS/AzCam interface changes)
- AGw1 - tested guide probes at different LUCI orientations. X-axis motions were ~15 seconds. Full field of AGW recommended for use.
- Updated spec and CI scripts for LASAIR and FACSUM
- monitor AGw1 currents and temps
- FACSUM radar image
- OVMS DSP code debug for Microgate
- OVMS new DSP code test on the telescope
- OVMS new interface document review
- low level AO SW for OVMS
- AdamHouseKeeper changes for ASM safety
- Monday morning off-office
- OVMS test on the telescope
- FW update and test?
- AdamHouseKeeper changes
- AdSec GUI for SOUL/FLAO warning
- 4D visit
- OC, WPC, Management
- sci-arc webinar - 21 attendees
- SHARK NIR/VIS s/w discussions
- DMS telem - evaluate chart.js/dygraph/plotly.js, needs polishing but getting there
- CH panel interview
- ENG1/2 time - launch a 64b w/s ?
- s/w group staff responsibilities and time allocation
- DMS ks, telem
- Follow-up with Olga on requirements of her LBC flats wiki page.
- Stephen re-kickstarted the ao-cravat VM, and some reorganization of the old "analysis" VM as well.
- Started on a major redesigning of the web interface, continue to work on it in the near future.
- Fixed a few issues with Chart.js plotting.
- Started to complete the database design.
- Merged the previous ApplyWaffleAO pull request.
- Focused on getting the new instrument authorization done in recent time.
- Working on entrace window dust detection algorithm for LUCI. Sent to Dave T. a bunch of profile observations for imaging and filters H, J or K. Although there is a signal is not clear how to make sense out of it.
- Revisiting Shark-Nir OT implementation.
- At Matthiue's request I looked through the git logs and and code for the OT and LUCI trying to find a elegible code to present to our software candidate.
- Worked on Arizona spreadsheet. I think I now understand what they want and how to do it.
- Still working on RTD issue #7642 . Near completion but I still need to run a few more tests.
- Possibly working on entrace window dust detection algorithm for LUCI
- If not then work on #7473, AFC pinhole issue
- Finish the Arizona spreadsheet
- Apply the AO/GUIDE star mix coordinate fix to production
- Microwave issues.
- rm507-2 unable to login Thursday night.
- re-install analysis.
- Cindy's computer to Windows 10
- VMs for archive folks
- Movie for DR.
- Pat is happy with WiFi.
- No entry on "summit slowdown". Haven't seen Jenny.
- broken Veeam, VM consolidation issues.
- Synology on mountain complaint Friday
- work on new obs, set a date with Steve A.
- re-racking machines in Tucson -- can we start moving before summer (DE)
- changes to dms kickstart, RPMs for some other NFS mounts.